# Model-free reinforcement learning with noisy actions for automated experimental control in optics

### v2

@@ Abstract
Setting up and controlling optical systems is often a challenging and tedious task. The high number of degrees of freedom to control mirrors, lenses, or phases of light makes automatic control challenging, especially when the complexity of the system cannot be adequately modeled due to noise or non-linearities. Here, we show that reinforcement learning (RL) can overcome these challenges when coupling laser light into an optical fiber, using a model-free RL approach that trains directly on the experiment without pre-training. By utilizing the sample-efficient algorithms Soft Actor-Critic (SAC) or Truncated Quantile Critics (TQC), our agent learns to couple with $90\%$ efficiency, comparable to the human expert. We demonstrate that direct training on an experiment can replace extensive system modeling. Our result exemplifies RL’s potential to tackle problems in optics, paving the way for more complex applications where full noise modeling is not feasible.
@@ 1 Introduction
In experimental physics, we work with complex and sensitive setups. Working in an optics lab means adjusting numerous mirrors, lenses, and other optical elements while optimizing complex parameters. Two of the main challenges are precision and the number of degrees of freedom. Often, tasks have to be repeated frequently. One example of such a task is coupling laser beams into optical fibers, used in many physics labs [1, 2, 3, 4]. It can be a laborious and time-consuming task, especially in experiments with many fibers. Automating tasks like this can, therefore, free up domain expertise for more challenging tasks. Most of these repeated tasks have a very clear goal and can either be described as alignment or control problems. Alignment means the correct steering of a laser beam through an optical setup. Control refers to maintaining a dynamic experiment at a desired position using feedback loops. While fiber coupling is primarily an alignment problem, correcting for drift can be considered control.
@@
Automation of alignment and control tasks is a classic use case of reinforcement learning (RL) [5, 6, 7]. RL has seen considerable success in recent years, both in general [8, 9, 10, 11, 12, 13, 14] and specifically in robotics [15, 16, 17, 18, 19]. However, due to many RL algorithms relying on a huge amount of data, at least in environments with continuous action spaces, most of these were performed in simulated or toy environments [20, 21]. Comparatively few experiments were done in real-world environments [22, 23, 24]. With the recent advance of more sample-efficient algorithms for continuous action spaces, like Deep Deterministic Policy Gradient (DDPG) [25], Twin Delayed Policy Gradient (TD3) [26], Soft Actor-Critic (SAC) [27], and Truncated Quantile Critics (TQC) [28], directly training in an experiment has become more feasible. However, we still face several challenges, such as partial observability, time-consuming training, and noise, when applying RL to real-world setups [22, 23].
@@
In this work, we demonstrate how an RL agent successfully learns to couple light into an optical fiber, reaching efficiencies comparable to those of a human expert. We set up an experiment for fiber coupling on an optical table, motorizing the mirrors that guide the laser beam into the fiber. Our goal is to reach a specific coupling efficiency, which is the fraction of light entering the fiber. We have not included the absolute motor positions in the observation to train our agent to improve the coupling efficiency for any type of misalignment. This means that the partial observability increases, but the agent generalizes to setup perturbations.
@@
The primary issue we encountered was the lack of precision in the motors. For instance, returning to a position was only possible with a considerable and unpredictable offset, which leads to noise in the actions, a special type of stochasticity of the environment. In contrast to adding artificial noise to the actions for exploration [29, 30, 31, 32, 33], in our case, the noisy actions are inherent to the system. To solve this problem without a full analysis and modeling of the behavior, we let our agent train directly on the experiment with the standard StableBaselines3 [34] implementations of SAC and TQC. To reset a training episode, we could not reliably move to an absolute position but implemented a reset procedure that mainly relied on relative movement steps.
@@
Despite the noisy actions and partial observability of the environment, the agent learns to reliably couple to an efficiency of $\geq 90\%\pm 2\%$ starting from a low power over the course of nearly four days. If we only need a smaller efficiency, e.g., $87\%\pm 2\%$, the training only takes twenty hours and can be performed in two nights, not taking away experimenting time. For comparison, the maximum coupling efficiency observed by the experimenter was $92\%$, and the one reached by the agent was $93\%$. We find that tuning the training parameters thoroughly is crucial to reducing training time, which is of high priority for real-world RL applications.
@@
Our successful training is a first step towards further applications. First, our experiment shows that laser beam alignment using RL is generally possible. A transfer to other scenarios, such as interference optimization of two beams at a beam splitter or alignment of a laser field to an optical resonator, is straightforward and requires only a change of sensor [35]. Secondly, with training directly on the experiment, we show an example of applying RL to control tasks without having to model the experiment in detail beforehand. This is particularly important for more complicated experiments, such as those in quantum and atomic optics, where it may be disproportionate or impossible to simulate the exact dynamics and noise.
@@ 2 Related work
The application of RL to optical systems encompasses a wide range of topics, including optical networks [36, 37, 38, 39, 40, 41, 42, 43, 44], adaptive optics [45, 46, 47, 48, 49, 50, 51, 52], and optical nanostructure, thin films and optical layers [53, 54, 55, 56]. More related to our problem is work that studies how RL can be used to align and control tabletop optical experiments with lasers. In this category, some works are realized merely on simulation. Examples include studying mode-locked lasers [57], combining laser beams [58], and stacking laser pulses [59, 60]. Other works include investigating how RL performs in an actual experiment. Most agents in these studies are, however, not trained in the experiment but in simulation. Examples include aligning an optical interferometer [61, 62], operating optical tweezers [63], combining laser beams [64] and operating pulsed lasers [65]. It is rare that the agent is trained directly on the experiment [24]. One example is a study combining pulsed laser beams [66]. Here, one actuator performs the actions, and the output is a scalar, the power. The training time is about 4 hours; simulations show that this would quickly go up to 1-2 days if more than two beams should be combined. Another example is the generation of a white light continuum [67]. Thereby, both the states and the actions are given by absolute positions of three actuators and moving to those positions, respectively. This gives their environment a relatively high observability for a real-world task. The authors claim to have obtained a successfully trained agent in 20 minutes. We deal with a higher (4) dimensional action space than both of these works. RL was also used to optimize the output power of an X-ray source [68]. For this, a single actuator was discretely controlled based on a scalar signal. As a side project, the paper looks at a simplified approach to fiber coupling using only two degrees of freedom and working with a discrete action space of size 4 employing DQNs [69]. The work does not present the achieved coupling efficiencies. In contrast, we work with continuous action spaces and control all four degrees of freedom necessary for general beam alignment required for optimal fiber coupling. While training on the experiment can be difficult for many reasons (see Section 3), it can be the last resort in cases where creating a model that accurately represents the noise and dynamics of the system is very time-consuming, if not infeasible. Using a too-inaccurate model, however, would make it impossible to cross the simulation-reality gap [70]. We therefore decided to study the little-explored field of in-situ training.
@@ 3.1 Experimental setup
To efficiently couple laser light into an optical fiber, we need a specific setup. Our goal is to reach a certain coupling efficiency, which is the fraction of light entering the fiber. To achieve this, the light has to enter the fiber at a specific angle and precise spot. The coupling efficiency depends on how accurately both are matched. To fulfill both constraints, we need two degrees of freedom in each axis, horizontal ($x$) and vertical ($y$[71, 72]. These are the same requirements as for any other beam alignment task. This means that two mirrors, each tiltable in $x$ and $y$, are sufficient to acquire an arbitrary beam alignment. Furthermore, the laser beam must have the correct size, which is achieved by placing lenses in the correct position before the light enters the fiber. To simplify the setup, we decided to motorize only the mirrors but not the lenses. In addition to the motorized mirrors, we have two mirrors that can be steered with hand-tuneable knobs. This makes it easier for humans to couple light into the fiber and can be used to introduce perturbations to the setup. We measure the power at the output of the fiber with a power meter (Thorlabs PM160). With the help of a reference power measurement, we can determine the coupling efficiency or normalized output power with an error of $2\%$. Our experiment is depicted in Figure 1, and further details are given in Appendix A.
@@ 3.2 RL challenges
When we use RL to fiber couple in our lab, we face several challenges. The training is time-consuming as one actuation step takes about 1 second. Furthermore, due to laser safety and possible equipment damage, we have to restrict the movement range of our actuators. The two challenges that are most crucial for shaping our environment are partial observability and motor imprecision.
@@ 4 Our method
We cannot write down a Markov state (see e.g. [5] for an introduction) for our system. Therefore, we treat it as an unknown episodic partially observable Markov decision process (POMDP, see e.g. [73]). When sampling from it, we get a stochastic process $o_{1},a_{1},r_{1},o_{2},a_{2},r_{2},...,o_{\tau}$, where $o_{t},a_{t}$ and $r_{t}$ are observations, actions and rewards at the discretized time $t$, and $\tau$ is the episode length limitited by the maximal episode length $T$, i.e. $\tau\leq T$. See Tables 1 and 2 for environment hyperparameters.
@@
We treat our action space as the 4-dimensional continuous action space $[-1,1]^{\times 4}$. At time $t$, we can decompose the action $a_{t}$ as $a_{t}=(a_{\mathrm{m}1x},a_{\mathrm{m}1y},a_{\mathrm{m}2x},a_{\mathrm{m}2y})$ where each component belongs to a different actuator. For example, $a_{\mathrm{m}1x}$ belongs to the actuator that tilts mirror $1$ in the horizontal ($x$) direction. Each of these actions is then multiplied by the maximum allowed action in actuator steps $a_{\mathrm{max}}$, rounded to the next integer, and sent to the different controllers. Using the virtual testbed, we find that maximum actions of approximately $a_{\mathrm{max}}=6\cdot 10^{3}$ are optimal (see Appendix C.6). However, in the experimental environment, the actuators have no feedback loop, so, potentially, they move significantly less due to their imprecision. This makes our actions noisy, which adds to the stochasticity in the environment. As discussed in Appendix C.8, this especially complicates the task for high powers.
@@
As we are dealing with a partially observable system, for successful training, we need to take great care in defining our observations. The only thing we observe in our environment is the coupling efficiency or normalized power, denoted $P\in[0,1]$. For example, $P_{t}$ denotes the normalized power at time $t$. As usual in POMDPs [66, 61]), we include a history of length $n\in\mathbb{N}$ in the observation. This would lead to an observation like $o_{t}=(P_{t-n},...,P_{t-1},P_{t})$ at time step $t$. It is common in RL experiments to observe the environment before and after an action, not during this action. However, this is not the only information available to us: In principle, we can record the power almost continuously while the actuators are moving, i.e., we can record $(P_{t},P_{t+1/m_{t}},...,P_{t+1})$ during action $a_{t}$ where $m_{t}$ is the number of powers measured during that time. In the virtual testbed, we noticed that it is beneficial to use some of this information in our observation (see Appendix C.5). In particular, we take the average power $P_{\mathrm{ave},t}=\sum_{i=0}^{m_{t}}P_{t-1+i/m_{t}}/(m_{t}+1)$ the maximal power observed $P_{\mathrm{max},t}=\max_{i=0,...,m_{t}}P_{t-1+i/m_{t}}$ and its relative position in the list of powers $x_{\mathrm{max},t}=(\mathrm{arg}\,\max_{i=0,...,m_{t}}P_{t-1+i/m_{t}})/(m_{t}+1)$ into account. In addition, it is helpful to use the performed actions as part of the observation. This leaves us with the observation
@@
We reset the environment either after a chosen maximum episode length $t=T\in\mathbb{N}$, when the agent reached its goal (i.e., $P_{t}>P_{\text{goal}}$), or the agent failed (i.e., $P_{t}<P_{\text{fail}}$). Using the virtual testbed, we find that it helps with reaching higher goals like $P_{\text{goal}}=0.9$ if we implement an instance of curriculum learning [74] by starting with lower goal powers and raising the goal power during training, especially starting from $P_{\text{start, goal}}=0.85$ (see Appendix C.4).
@@
The values of $0.09$ and ten episodes were determined empirically by observing the algorithm performing on the experiment. Before starting the episode, we always perform some random steps to randomize the process more. Our reset procedure introduces a small dependence between successive episodes. However, this did not affect the training performance in the virtual testbed much (see. Appendix C.3). Furthermore, due to the motors’ inaccuracies, full independence was not possible in this experiment.
@@
$=$ $\left\{\begin{array}[]{ll}-A_{f}\left((1-\alpha_{f})\exp{(-\beta_{f,1}\frac{t}{T})}+\alpha_{f}\exp{(-\beta_{f,2}\frac{P_{t}}{P_{\text{fail}}})}\right)&\text{if }P_{t}<P_{\text{fail}}\\ A_{g}\left((1-\alpha_{g})\exp{(-\beta_{g,1}\frac{t}{T})}+\alpha_{g}\exp{(\beta_{g,2}\frac{P_{t}}{P_{\mathrm{goal}}})}\right)&\text{if }P_{t}>P_{\text{goal}}\\ \frac{A_{s}}{T}\left((1-\alpha_{s})\exp{(\beta_{s}(P_{t}-P_{\mathrm{goal}}))}+\alpha_{s}\left(P_{t}-P_{\mathrm{min}}\right)\right)&\text{else }\end{array}\right.$
@@
The continuous action space limits our choice of algorithm. Of the algorithms tested in the virtual testbed (TQC, TD3, SAC, DDPG, PPO, and Advantage Actor-Critic (A2C)), PPO and A2C performed worst. As they both do not use a replay buffer, this was expected. They are closely followed by DDPG. SAC, TD3, and TQC performed much better. SAC performed almost always slightly better than TD3. TQC has a small drop compared to SAC in the middle of training that gets larger with rising goal powers. See Appendix C.7 for a discussion. We used the algorithms from StableBaselines3 [34] with standard hyperparameters further discussed in Appendix E. In the main experiment, we tested both TQC and SAC as algorithms.
For both the main experiment and virtual testbed, we used a gymnasium environment [75] and the parameters provided in Tables 1 and 2 if not specified differently. Our strategy is that the agent learns to deal with the noisy actions directly in the experiment. We are especially interested in investigating this in-situ learning of noise as, in our area, we are often dealing with noise sources that cannot be modeled accurately, for example, when dealing with quantum states of light. In these cases, the only solution will be to learn to handle the noise through direct interaction with the experiment.

### v3

@@ Abstract
Setting up and controlling optical systems is often a challenging and tedious task. The high number of degrees of freedom to control mirrors, lenses, or phases of light makes automatic control challenging, especially when the complexity of the system cannot be adequately modeled due to noise or non-linearities. Here, we show that reinforcement learning (RL) can overcome these challenges when coupling laser light into an optical fiber, using a model-free RL approach that trains directly on the experiment without pre-training on simulations. By utilizing the sample-efficient algorithms Soft Actor-Critic (SAC), Truncated Quantile Critics (TQC), or CrossQ, our agents learn to couple with $90\%$ efficiency. A human expert reaches this efficiency, but the RL agents are quicker. In particular, the CrossQ agent outperforms the other agents in coupling speed while requiring only half the training time. We demonstrate that direct training on an experiment can replace extensive system modeling. Our result exemplifies RL’s potential to tackle problems in optics, paving the way for more complex applications where full noise modeling is not feasible.
@@ 1 Introduction
In experimental physics, we work with complex and sensitive setups. Working in an optics lab means adjusting numerous mirrors, lenses, and other optical elements while optimizing complex parameters. Two of the main challenges are precision and the number of degrees of freedom. Often, tasks have to be repeated frequently. One example of such a task is coupling laser beams into optical fibers, used in many physics labs (Addanki et al., 2018; Lu et al., 2019; Larsen et al., 2019; Xavier & Lima, 2020). It can be a laborious and time-consuming task, especially in experiments with many fibers. Automating tasks like this can, therefore, free up domain expertise for more challenging tasks. Most of these repeated tasks have a very clear goal and can either be described as alignment or control problems. Alignment means the correct steering of a laser beam through an optical setup. Control refers to maintaining a dynamic experiment at a desired position using feedback loops. While fiber coupling is primarily an alignment problem, correcting for drift can be considered control.
@@
Automation of alignment and control tasks is a classic use case of reinforcement learning (RL) (Sutton & Barto, 2018; Krenn et al., 2023; Vernuccio et al., 2022). RL has seen considerable success in recent years, both in general (Silver et al., 2017; 2018; Mnih et al., 2015; Vinyals et al., 2019; Berner et al., 2019; Fawzi et al., 2022; Ruiz et al., 2024) and specifically in robotics (James et al., 2019; Andrychowicz et al., 2020; Kumar et al., 2023; Sontakke et al., 2023; Zhao et al., 2020). However, due to many RL algorithms relying on a huge amount of data, at least in environments with continuous action spaces, most of these were performed in simulated or toy environments (Akkaya et al., 2019; Ranaweera & Mahmoud, 2023). Comparatively few experiments were done in real-world environments (Dulac-Arnold et al., 2019; 2021; Ding & Dong, 2020). With the recent advance of more sample-efficient algorithms for continuous action spaces, like Deep Deterministic Policy Gradient (DDPG) (Lillicrap et al., 2019), Twin Delayed Policy Gradient (TD3) (Fujimoto et al., 2018), Soft Actor-Critic (SAC) (Haarnoja et al., 2018), Truncated Quantile Critics (TQC) (Kuznetsov et al., 2020), and CrossQ (Bhatt et al., 2024) directly training in an experiment has become more feasible. However, we still face several challenges, such as partial observability, time-consuming training, and noise, when applying RL to real-world setups (Dulac-Arnold et al., 2019; 2021).
@@
In this work, we demonstrate how an RL agent successfully learns to couple light into an optical fiber, reaching efficiencies equivalent to those of a human expert. We set up an experiment for fiber coupling on an optical table, motorizing the mirrors that guide the laser beam into the fiber. Our goal is to reach a specific coupling efficiency, which is the fraction of light entering the fiber. We have not included the absolute motor positions in the observation to train our agent to improve the coupling efficiency for any type of misalignment. This means that the partial observability increases, but the agent generalizes to setup perturbations.
@@
The primary issue we encountered was the lack of precision in the motors. For instance, returning to a position was only possible with a considerable and unpredictable offset, which leads to noise in the actions, a special type of stochasticity of the environment. In contrast to adding artificial noise to the actions for exploration (Chiappa et al., 2023; Eberhard et al., 2022; Hollenstein et al., 2022; Zhang & Hoof, 2021; Plappert et al., 2018), in our case, the noisy actions are inherent to the system. To solve this problem without a full analysis and modeling of the behavior, we let our agent train directly on the experiment with the standard StableBaselines3 (Raffin et al., 2021) implementations of SAC, TQC, and CrossQ. To reset a training episode, we could not reliably move to an absolute position but implemented a reset procedure that mainly relied on relative movement steps.
@@
Despite the noisy actions and partial observability of the environment, the agent learns to reliably couple to an efficiency of $\geq 90\%\pm 2\%$ starting from a low power over the course of two days using CrossQ or SAC and nearly four days when trained with TQC. If we only need a smaller efficiency, e.g., $87\%\pm 2\%$, the training only takes twenty hours and can be performed in two nights, not taking away experimenting time. For comparison, the maximum coupling efficiency observed by the experimenter was $92\%$, and the one reached by the agent was $93\%$. We find that tuning the environment parameters thoroughly is crucial to reducing training time, which is of high priority for real-world RL applications.
@@
Our successful training is a first step towards further applications. First, our experiment shows that laser beam alignment using RL is generally possible. A transfer to other scenarios, such as interference optimization of two beams at a beam splitter or alignment of a laser field to an optical resonator, is straightforward and requires only a change of sensor (Siegman, 1986). Secondly, with training directly on the experiment, we show an example of applying RL to control tasks without having to model the experiment in detail beforehand. This is particularly important for more complicated experiments, such as those in quantum and atomic optics, where it may be disproportionate or impossible to simulate the exact dynamics and noise.
@@ 2 Related work
Reinforcement learning is widely used in robotics (Singh et al., 2022). In many real-world robotics environments, the challenge of noisy observations is apparent, while there is less focus on noisy or inaccurate actions (Zhao et al., 2020). Still, in sim-to-real transfer studies, different forms of actuator imprecision have been shown to significantly affect policy transferability, but agents can still learn in environments with backlash (Golemo et al., 2018). We identified backlash as a key challenge for our agent, rather than the widely studied problem of noisy observations. As the robotics field is highly diverse and direct comparisons between experiments are generally difficult, we focus our discussion of related work on the subfield of optical experiments.
@@
The application of RL to optical systems encompasses a wide range of topics, including optical networks (Kiran et al., 2007; Suárez-Varela et al., 2019; Chen et al., 2019b; a; Luo et al., 2019; Troia et al., 2019; Li et al., 2020; Natalino & Monti, 2020; Yu et al., 2023), adaptive optics (Ke et al., 2019; Nousiainen et al., 2021; Durech et al., 2021; Landman et al., 2021; Nousiainen et al., 2022a; b; Pou et al., 2022; Nousiainen et al., 2024), and optical nanostructure, thin films and optical layers (Sajedian et al., 2019; Jiang et al., 2020; Wang et al., 2021; Wankerl et al., 2021). More related to our problem is work that studies how RL can be used to align and control tabletop optical experiments with lasers. In this category, some works are realized merely on simulation. Examples include studying mode-locked lasers (Sun et al., 2020), combining laser beams (Tünnermann & Shirakawa, 2021), and stacking laser pulses (Abuduweili et al., 2021; Abuduweili & Liu, 2022). Other works include investigating how RL performs in an actual experiment. Most agents in these studies are, however, not trained in the experiment but in simulation. Examples include aligning an optical interferometer (Sorokin et al., 2020; Mukund et al., 2023), operating optical tweezers (Praeger et al., 2021), combining laser beams (Shpakovych et al., 2021) and operating pulsed lasers (Kuprikov et al., 2022). It is rare that the agent is trained directly on the experiment (Ding & Dong, 2020). One example is a study combining pulsed laser beams (Tünnermann & Shirakawa, 2019). Here, one actuator performs the actions, and the output is a scalar, the power. The training time is about 4 hours; simulations show that this would quickly go up to 1-2 days if more than two beams should be combined. Another example is the generation of a white light continuum (Valensise et al., 2021). Thereby, both the states and the actions are given by absolute positions of three actuators and moving to those positions, respectively. This gives their environment a relatively high observability for a real-world task. The authors claim to have obtained a successfully trained agent in 20 minutes. We deal with a higher (4) dimensional action space than both of these works. RL was also used to optimize the output power of an X-ray source (Mareev et al., 2023). For this, a single actuator was discretely controlled based on a scalar signal. As a side project, the paper looks at a simplified approach to fiber coupling using only two degrees of freedom and working with a discrete action space of size 4 employing DQNs (Mnih et al., 2013). The work does not present the achieved coupling efficiencies. In contrast, we work with continuous action spaces and control all four degrees of freedom necessary for general beam alignment required for optimal fiber coupling. While training on the experiment can be difficult for many reasons (see Section 3), it can be the last resort in cases where creating a model that accurately represents the noise and dynamics of the system is very time-consuming, if not infeasible. Using a too-inaccurate model, however, would make it impossible to cross the simulation-reality gap (Salvato et al., 2021). We therefore decided to study the little-explored field of in-situ training.
@@ 3.1 Experimental setup
To efficiently couple laser light into an optical fiber, we need a specific setup. Our goal is to reach a certain coupling efficiency, which is the fraction of light entering the fiber. To achieve this, the light has to enter the fiber at a specific angle and precise spot. The coupling efficiency depends on how accurately both are matched. To fulfill both constraints, we need two degrees of freedom in each axis, horizontal ($x$) and vertical (yy(Senior, 2009; Knigge et al., 2021). These are the same requirements as for any other beam alignment task. This means that two mirrors, each tiltable in $x$ and $y$, are sufficient to acquire an arbitrary beam alignment. Furthermore, the laser beam must have the correct size, which is achieved by placing lenses in the correct position before the light enters the fiber. To simplify the setup, we decided to motorize only the mirrors but not the lenses. In addition to the motorized mirrors, we have two mirrors that can be steered with hand-tuneable knobs. This makes it easier for humans to couple light into the fiber and can be used to introduce perturbations to the setup. We measure the power at the output of the fiber with a power meter (Thorlabs PM160). With the help of a reference power measurement, we can track power fluctuations and determine the coupling efficiency or normalized output power with an error of $2\%$. Our experiment is depicted in Figure 1, and further details are given in Appendix A.
@@ 3.2 RL challenges
When we use RL to fiber couple in our lab, we face several challenges. The training is time-consuming as one environment step takes about one second. Furthermore, due to laser safety and possible equipment damage, we have to restrict the movement range of our actuators. The two challenges that are most crucial for shaping our environment are partial observability and motor imprecision.
@@ 4 Our method
We cannot write down a Markov state (see e.g. Sutton & Barto (2018) for an introduction) for our system. Therefore, we treat it as an unknown episodic partially observable Markov decision process (POMDP, see e.g. Graesser & Keng (2020)). When sampling from it, we get a stochastic process $o_{1},a_{1},r_{1},o_{2},a_{2},r_{2},...,o_{\tau}$, where $o_{t},a_{t}$ and $r_{t}$ are observations, actions and rewards at the discretized time $t$, and $\tau$ is the episode length limitited by the maximal episode length $T$, i.e. $\tau\leq T$. See Tables 1 and 2 for environment hyperparameters.
@@
We treat our action space as the 4-dimensional continuous action space $1,1]^{\times 4}$. At time $t$, we can decompose the action $a_{t}$ as $a_{t}=(a_{\mathrm{m}1x},a_{\mathrm{m}1y},a_{\mathrm{m}2x},a_{\mathrm{m}2y})$ where each component belongs to a different actuator. For example, $a_{\mathrm{m}1x}$ belongs to the actuator that tilts mirror $1$ in the horizontal ($x$) direction. Each of these actions is then multiplied by the maximum allowed action in actuator steps $a_{\mathrm{max}}$, rounded to the next integer, and sent to the different controllers. Using the virtual testbed, we find that maximum actions of approximately $a_{\mathrm{max}}=6\cdot 10^{3}$ are optimal (see Appendix C.6). However, in the experimental environment, the actuators have no feedback loop, so, potentially, they move significantly less due to their imprecision. This makes our actions noisy, which adds to the stochasticity in the environment. As discussed in Appendix C.8, this especially complicates the task for high powers.
@@
As we are dealing with a partially observable system, for successful training, we need to take great care in defining our observations. The only thing we observe in our environment is the coupling efficiency or normalized power, denoted $P\in[0,1]$. For example, $P_{t}$ denotes the normalized power at time $t$. As usual in POMDPs (Tünnermann & Shirakawa, 2019; Sorokin et al., 2020)), we include a history of length $n\in\mathbb{N}$ in the observation. This would lead to an observation like $o_{t}=(P_{t-n},...,P_{t-1},P_{t})$ at time step $t$. It is common in RL experiments to observe the environment before and after an action, not during this action. However, this is not the only information available to us: In principle, we can record the power almost continuously while the actuators are moving, i.e., we can record $(P_{t},P_{t+1/m_{t}},...,P_{t+1})$ during action $a_{t}$ where $m_{t}$ is the number of powers measured during that time. In the virtual testbed, we noticed that it is beneficial to use some of this information in our observation (see Appendix C.5). In particular, we take the average power $P_{\mathrm{ave},t}=\sum_{i=0}^{m_{t}}P_{t-1+i/m_{t}}/(m_{t}+1)$ the maximal power observed $P_{\mathrm{max},t}=\max_{i=0,...,m_{t}}P_{t-1+i/m_{t}}$ and its relative position in the list of powers $x_{\mathrm{max},t}=(\mathrm{arg}\,\max_{i=0,...,m_{t}}P_{t-1+i/m_{t}})/(m_{t}+1)$ into account. In addition, it is helpful to use the performed actions as part of the observation. This leaves us with the observation
@@
We reset the environment either after a chosen maximum episode length $t=T\in\mathbb{N}$, when the agent reached its goal (i.e., $P_{t}>P_{\text{goal}}$), or the agent failed (i.e., $P_{t}<P_{\text{fail}}$). While training with TQC on the virtual testbed, we find that it helps with reaching higher goals like $P_{\text{goal}}=0.9$ if we implement an instance of curriculum learning (Narvekar et al., 2020) by starting with lower goal powers and raising the goal power during training, especially starting from $P_{\text{start, goal}}=0.85$ (see Appendix C.4).
@@
The values of $0.09$ and ten episodes were determined empirically by observing the algorithm performing on the experiment. Before starting the episode, we always perform some random steps to randomize the process more. Our reset procedure introduces a small dependence between successive episodes. Full independence was not possible in this experiment due to the motors’ inherent inaccuracies. In contrast, a fully random-based reset procedure is applicable while training solely on the virtual testbed. We compared this with our custom reset method on the virtual testbed and found no major effect on the training performance (see Appendix C.3).
@@
[…]
@@
The continuous action space limits our choice of algorithm. Our main environment parameter searches on the virtual testbed were performed with TQC. Subsequently, we also tested CrossQ, TD3, SAC, DDPG, PPO, and Advantage Actor-Critic (A2C). Of these, PPO and A2C performed worst. As they both do not use a replay buffer, this was expected. They are closely followed by DDPG. CrossQ, SAC, TD3, and TQC performed much better. SAC performed almost always slightly better than TD3. TQC has a small drop compared to SAC in the middle of training that gets larger with rising goal powers. With rising goal powers CrossQ shows an advantage over the other algorithms. See Appendix C.7 for a discussion. Thus, in the main experiment, we tested CrossQ, TQC, and SAC as algorithms. We used the algorithms from StableBaselines3 (Raffin et al., 2021) with standard hyperparameters further discussed in Appendix E.
Comparability of the virtual testbed and the real-world experiment
The experiment has two main differences from the virtual testbed. Firstly, the actions are highly noisy due to the motors’ inaccuracies. Secondly, as described in Section 3, the two elements of the action belonging to one axis are coupled. However, in the virtual testbed, to keep the design simple, we uncoupled these actions. The interaction of the different actuators to achieve high coupling efficiencies is therefore easier in the virtual testbed, and for the agent, it is harder to learn a working policy in the real experiment. The ablation studies on the observation done on the virtual testbed (see Appendix C.5) are generally transferable to the real-world experiment: as the level of partial observability is similar in both environments and the dimension of the action space is the same the studies of the history length are fully transferable (see Figure 10 a). The observation design (see Figure 10 b) is also transferable, as we have low noise in the power sensors, and saving these values is the same in both environments. Given that the power is the only environmental value utilized in the reward, and considering its low noise, the reward design is likewise transferable.
For both the main experiment and virtual testbed, we used a gymnasium environment (Brockman et al., 2016) and the parameters provided in Tables 1 and 2 if not specified differently. Our strategy is that the agent learns to deal with the noisy actions directly in the experiment. We are especially interested in investigating this in-situ learning of noise as, in our area, we are often dealing with noise sources that cannot be modeled accurately, for example, when dealing with quantum states of light. In these cases, the only solution will be to learn to handle the noise through direct interaction with the experiment.